Bill is 29 yrs older than his daughter Rebecca. In 15 years time their ages will be as such that Bill will be twice as old as rebecca. How old is Bill now?( Let bill's age be x years)
I tried
x + 15 = 3 ( x - 29 )

Answer by Paul(988)

Rebecca = x
Bill = x+29

In fiftheen years there ages:
x+29+15
and
x+15

2(x+15)=x+29+15
2x+30=x+44
x=14

14+29=43

Hence, Rebecca is 14 years old and Bill is 43 years old.
PAUL.